You are given a small pipe with a diameter of 1 mm flowing
cooling fluid (water) at a temperature of 20C. Fluid flow velocity is regulated at a constant speed
of 0.2 m / s and assumed to be uniform. If the pipe surface is given a constant heat flux
of 6000 W / m2
and you assume the analysis is done in a fully develop area, then
based on your design calculations, at what length of the pipe (measured from the beginning of the input)
water temperature will increase to 74C? Note: the value of thermal diffusivity and viscosity
the kinematic is 1,541 ? 10−7 m^2s and 0.556 ? 10−6 m^2s2. Ketika anda kerja praktek di industri oil and gas, kemungkinan besar anda akan menjumpai banyak sekali kasus yang melibatk, ? = 0.6367 W / mk
Answer #1

Assumption -

Fully develop region.

Calculate -
Length of the pipe at which water temperature will increase to 74C

Given data -

v (Kinematic viscosity) = 0.556 ? 10−6 m^2s

? = 0.6367 W / mk

Gooow/m2 To Ta 200 14 w

First, calculate whether the flow is laminar or turbulent -

Re=\frac{vL}{\nu }

Where

v = velocity of fluid i.e. water

\nu = Kinematic viscosity m2/sec

L = Hydrodynamic or characteristic length i.e. the length at which the flow is in contact.

Re=\frac{0.2\times 0.001}{0.556 \times 10^{-6} }

Re= 359.7122 < 2000

For pipe flow, boundary layer growth is almost similar to the flow over a flat plate.

Now using empirical relation of average nusselt number for laminar flow i.e.

Nu_x = 0.332(Re_x)^{0.5}(Pr)^{0.33}

Where

Nu = Nusselt number

Re = Reynold number

Pr = Prandtl Number

\frac{hL}{k}= 0.664(Re)^{0.5}(Pr)^{0.33}

Where

h = Convection heat transfer coefficient

L = D(For cylinder) =Hydrodynamic or characteristic length

k = Thermal conductivity in W/m-K

Pr = \frac{\mu C_p}{k}

Pr = \frac{\nu }{\alpha }=\frac{\mu C_p}{k}

Where

Kinematic viscosity

0.556\times 10^{-6}= \frac{\mu}{1000}

Therefore,

Dynamic viscosity

\mu=0.556\times 10^{-3}Pa-sec

Specific heat of water = 4.18 KJ/Kg-K

? = 0.6367 W / mk

Pr = \frac{0.556\times 10^{-3}\times 4.18\times 1000}{0.6367}

Pr = 3.65

or

Pr = \frac{\nu }{\alpha }=3.6119 (Use this one because both values are given in question directly)

Now assume 74 C temperature occurs at distance x from the leading edge.

Therefore the Reynold number will be -

Re=\frac{0.2\times x}{0.556 \times 10^{-6} }

Now assume 74 C temperature occurs at distance x from the leading edge.

Therefore,

Put the above values in empirical relation -

h_{local}= 0.332\times \frac{0.6367}{x}(359712.2x)^{0.5}(3.6119)^{0.33}

h_{x}= \frac{193.6883}{\sqrt{x}}

See at the surface of the pipe, velocity is zero or approximately zero. So heat flux at the surface is given as -

\frac{Q}{A}=h_{x}(T_s-T_w)

6000=\frac{193.6883}{\sqrt{x}}(74-20)

\mathbf{x= 3.038 m}

If you have any doubt or answer does not match please write in comment section.

Thank you.